# Recent Trends of Multimodal Affective Computing: A Survey from NLP Perspective

### v1

@@ Abstract
Multimodal affective computing (MAC) has garnered increasing attention due to its broad applications in analyzing human behaviors and intentions, especially in text-dominated multimodal affective computing field. This survey presents the recent trends of multimodal affective computing from NLP perspective through four hot tasks: multimodal sentiment analysis, multimodal emotion recognition in conversation, multimodal aspect-based sentiment analysis and multimodal multi-label emotion recognition. The goal of this survey is to explore the current landscape of multimodal affective research, identify development trends, and highlight the similarities and differences across various tasks, offering a comprehensive report on the recent progress in multimodal affective computing from an NLP perspective. This survey covers the formalization of tasks, provides an overview of relevant works, describes benchmark datasets, and details the evaluation metrics for each task. Additionally, it briefly discusses research in multimodal affective computing involving facial expressions, acoustic signals, physiological signals, and emotion causes. Additionally, we discuss the technical approaches, challenges, and future directions in multimodal affective computing. To support further research, we released a repository that compiles related works in multimodal affective computing, providing detailed resources and references for the community  [^1] https://github.com/LeMei/Affective-Computing.
@@ I Introduction
Affective computing combines expertise in computer science, psychology, and cognitive science and its goal is to equip machines with the ability to recognize, interpret, and emulate human emotions [1, 2, 3, 4, 5, 6]. Today, the world around us comprises various modalities—we perceive objects visually, hear sounds audibly, feel textures tangibly, smell odors olfactorily, and so forth. A modality refers to the way an experience is perceived or occurs, and is often associated with sensory modalities such as vision or touch, which are essential for communication and sensation. The significant advancements in multimodal learning across various fields [7, 8] garnered increasing attention and accelerated the progress of multimodal affective computing.
@@
Multimodal affective computing seeks to develop models capable of interpreting and reasoning sentiment or emotional state over multiple modalities. In its early stages, researchers of affective computing predominantly focused on unimodal tasks, examining text-based, audio-based, and vision-based affective computing separately. For instance, D-MILN [9] is a textual sentiment classification model, while work [10] utilizes BiLSTM models trained on raw audio to predict the average sentiment of crowd responses. Today, sentiment analysis is widely employed across various modalities for applications such as market research, brand monitoring, customer service analysis, and social media monitoring. Recent advancements in multimedia technology [11, 12, 13, 14] have diversified the channels for information dissemination, with an influx of news, social media platforms like Weibo, and video content. These developments have integrated textual (spoken features), acoustic (rhythm, pitch), and visual (facial attributes) information to comprehensively analyze human emotions. For example, Xu et al.  [15] introduces image modality data into traditional text-based aspect-level sentiment analysis, creating the new task of multimodal aspect-based sentiment analysis. Similarly, Wang et al. [16] extends textual emotion-cause pair extraction to a multimodal conversation setting, utilizing multimodal signals (text, audio, and video) to enhance the model’s ability to understand the emotions and their causes.
@@
Multimodal affective computing tasks closely related to several learning paradigms in machine learning, including transfer learning [17, 18, 19], multimodal learning [20, 21], muti-task learning  [22, 23, 24] and semantic understanding [25, 26]. Regarding transfer learning, it allows affective analysis models trained in one domain to be adapted for effective performance in different domains. By fine-tuning pre-trained models on limited data from the target domain, these models can be transferred to new domain, thereby enhancing their performance in multimodal affective computing tasks. In multimodal learning, cross-modal attention dynamically aligns and focuses on relevant information from different modalities, enhancing the model’s ability to capture sentiment by highlighting key features and their interactions. In multi-task learning, shared representations across affective computing tasks and modalities improve performance by capturing common sentiment-related features from text, audio, and video.
@@
More recently, the studies of multimodal learning have advanced the field by pre-training multimodal models on extensive multimodal datasets over years, further improving the performance on downstream tasks such as multimodal sentiment analysis [27, 28, 29, 30]. With the scaling of the pre-trained model, parameter-efficient transfer learning emerges such as adapter [31], prompt [32], instruction-tuning [33] and in-context learning [34, 35]. More and more works of multimodal affective computing leverage these parameter-efficient transfer learning methods to transfer knowledge from pre-trained models (e.g., unimodal pre-trained model or multimodal pre-trained model) to downstream affective tasks to improve model performance by further fine-tuning the pre-trained model. For instance, Zou et al. [36] design a multimodal prompt Transformer (MPT) to perform cross-modal information fusion. UniMSE [37] proposes an adapter-based modal fusion method, which injects acoustic and visual signals into the T5 model to fuse them with multi-level textual information.
@@
Multimodal affective computing encompasses tasks like sentiment analysis, opinion mining, and emotion recognition using modalities such as text, audio, images, video, physiological signals, and haptic feedback. This survey focuses mainly on three key modalities: natural language, visual signals, and vocal signals. We highlight four main tasks in this survey: Multimodal Sentiment Analysis (MSA), Multimodal Emotion Recognition in Conversation (MERC), Multimodal Aspect-Based Sentiment Analysis (MABSA), and Multimodal Multilabel Emotion Recognition (MMER). A considerable volume of studies exists in the field of multimodal affective computing, and several reviews have been published [38, 39, 40, 14]. However, these reviews primarily focus on specific affective computing tasks or specific single modality and overlook an overview of multimodal affective computing across multiple tasks, and the consistencies and differences among these tasks. The goal of this survey is twofold. First, this survey aims to provide a comprehensive overview of multimodal affective computing for beginners exploring deep learning in emotion analysis, detailing tasks, inputs, outputs, and relevant datasets. Second, it also offers insights for researchers to reflect on past developments, explore future trends, and examine technical approaches, challenges, and research directions in areas such as multimodal sentiment analysis and emotion recognition.
@@ IV Modal Feature Extractor
For audio modality, raw acoustic input needs to be processed into numerical sequential vectors. The common way is to use librosa  [^2] https://github.com/librosa/librosa. to extract Mel-spectrogram as audio features. It is the short-term power spectrum of sound and is widely used in modern audio processing. Transformer structure has achieved tremendous success of in the field of natural language processing and computer vision. Gong et al. [61] propose audio spectrogram Transformer (AST), which converted waveform into a sequence of 128-dimensional log Mel filterbank (fbank) features to encode audio modality.
@@ VI-B Pre-trained Model
In recent years, large language model (LLM) [56, 210] and multimodal pre-trained model [211, 26, 21, 212] has achieved significant progress [25, 210, 213]. Compared with non-pretrained model, pre-trained model contains massive transferred knowledge [31, 27], which can be introduced into multimodal represention learning to probe the richer information. Fig. 6 shows the use of pre-trained model in multimodal affective learning task.
@@ VI-C Enhanced Knowledge
External knowledge in machine learning and AI refers to information from outside the training dataset, including knowledge bases, text corpora, knowledge graphs, pre-trained models, and expert insights. Integrating this knowledge can improve performance, generalization, interpretability, and robustness to noisy or limited data. Fig. 7 shows the common way of incorporating external knowledge into multimodal affective learning task.
@@ VI-D4 Multimodal Multi-label Emotion Recognition
MERC MELD [239] T,A,V Friends TV English 7,532
IEMOCAP [240] T,A,V Act English 13,707
@@ VII-B Multimodal Emotion Recognition in Conversation
  • IEMOCAP [240] has 13,707 video clips of multi-party conversations, with labels following Ekman’s six universal emotions, including joy, sadness, fear, anger, surprise and disgust.
@@ IX-A Other Multimodal Affective Computing
Facial expression recognition has significantly evolved over the years, progressing from static to dynamic methods. Initially, static facial expression recognition (SFER) relied on single-frame images, utilizing traditional image processing techniques such as Local Binary Patterns (LBP) and Gabor filters to extract features for classification. The advent of deep learning brought Convolutional Neural Networks (CNNs), which markedly improved the accuracy of SFER [260, 261, 262, 263]. However, static methods were limited in capturing the temporal dynamics of facial expressions [264]. Some methods attempt to approach the problem from a local-global feature perspective, extracting more fine-grained visual representations and identifying key informative segments [265, 266, 267, 268, 269, 270]. These approaches enhance robustness against noisy frames, enabling uncertainty-aware inference. To further enhance accuracy, recent advancements in DFER focus on integrating multimodal data and employing parameter-efficient fine-tuning (PEFT) to adapt large pre-trained models for enhanced performance. Liu et al. [271] introduces the concept of expression reenactment (i.e. normalization), harnessing generative AI to mitigate noise in in-the-wild datasets. Moreover, the burgeoning evidential deep learning (EDL) has shown considerable promise by enabling explicit uncertainty quantification through the distributional measurement in latent spaces for improved interpretability, with demonstrated efficacy in zero-shot learning [272], multi-view classification [273, 274, 275], video understanding [276, 277, 278] and multi-modal named entity recognition.
@@ X Future work
Natural language (spoken and written), visual data (images and videos), and auditory signals (speech, sound, and music) have long been central to multimodal affective computing. Recently, new sensing data types, like haptic and ECG signals, are gaining attention. Haptic signals, which involve touch and convey sensory and emotional attributes, enhance user experiences in areas such as gaming, virtual reality, and mobile apps [317]. These signals offer immediate feedback and can improve user engagement. As research progresses, less-studied modalities like haptic will likely become crucial, complementing established methods and advancing the field of affective computing.
@@ XI Conclusion
Multimodal Affective Computing (MAC) has emerged as a crucial research direction in artificial intelligence with significant progress in understanding and interpreting emotions. This survey provides a comprehensive overview of the diverse tasks associated with multimodal affective computing, covering its research background, definitions, related work, technical approaches, benchmark dataset, and evaluation metrics. We group multimodal affective computing across MSA, MERC, MABSA and MMER tasks into four categories: multi-task learning, pre-trained modal, enhanced knowledge and context information. Additionally, we summarize the consistency and differences among various affective computing tasks. Also, we report the inherent challenges in multimodal sentiment analysis and explore potential directions for future research and development.

### v2

# Recent Trends of Multimodal Affective Computing: A Survey from an NLP Perspective

@@ Abstract
Multimodal affective computing has garnered increasing attention due to its broad applications in analyzing human behaviors and intentions, especially in text-dominated multimodal affective computing field. This survey presents the recent trends of multimodal affective computing from an NLP perspective through four hot tasks: multimodal sentiment analysis, multimodal emotion recognition in conversation, multimodal aspect-based sentiment analysis and multimodal multi-label emotion recognition. The goal of this survey is to explore the current landscape of multimodal affective research, identify development trends, and highlight the similarities and differences across various tasks, offering a comprehensive report on the recent progress in multimodal affective computing from an NLP perspective. This survey covers the formalization of tasks, provides an overview of relevant works, describes benchmark datasets, and details the evaluation metrics for each task. Additionally, it briefly discusses research in multimodal affective computing involving facial expressions, acoustic signals, physiological signals, and emotion causes. Additionally, we discuss the technical approaches, challenges, and future directions in multimodal affective computing. Finally, we release a repository that compiles related works in multimodal affective computing, providing detailed resources and references for the community [^1] https://github.com/LeMei/Affective-Computing.
@@ I Introduction
Affective computing combines expertise in computer science, psychology, and cognitive science and its goal is to equip machines with the ability to recognize, interpret, and emulate human emotions [1, 2, 3, 4, 5, 6, 7]. Today, the world around us comprises various modalities—we perceive objects visually, hear sounds audibly, feel textures tangibly, smell odors olfactorily, and so forth. A modality refers to the way an experience is perceived or occurs, and is often associated with sensory modalities such as vision or touch, which are essential for communication and sensation. The significant advancements in multimodal learning across various fields [8, 9] garnered increasing attention and accelerated the progress of multimodal affective computing. Multimodal affective computing seeks to develop models capable of interpreting and reasoning sentiment or emotional state over multiple modalities.
@@
In its early stages, researchers of affective computing predominantly focused on unimodal tasks, examining text-based, audio-based, and vision-based affective computing separately. For instance, D-MILN [10] is a textual sentiment classification model, while work [11] utilizes BiLSTM models trained on raw audio to predict the average sentiment of crowd responses. Today, sentiment analysis is widely employed across various modalities for applications such as market research, brand monitoring, customer service analysis, and social media monitoring. Recent advancements in multimedia technology [12, 13, 14, 15] have diversified the channels for information dissemination, with an influx of news, social media platforms like Weibo, and video content. These developments have integrated textual (spoken features), acoustic (rhythm, pitch), and visual (facial attributes) information to comprehensively analyze human emotions. For example, Xu et al. [16] introduced image modality data into traditional text-based aspect-level sentiment analysis, creating the new task of multimodal aspect-based sentiment analysis. Similarly, Wang et al. [17] extended textual emotion-cause pair extraction to a multimodal conversation setting, utilizing multimodal signals to enhance the model’s ability to understand the emotions and their causes.
@@
Multimodal affective computing tasks closely related to several learning paradigms in machine learning, including transfer learning [18, 19, 20], multimodal learning [21, 22], muti-task learning [23, 24, 25] and semantic understanding [26, 27]. Regarding transfer learning, it allows affective analysis models trained in one domain to be adapted for effective performance in different domains. By fine-tuning pre-trained models on limited data from the target domain, these models can be transferred to new domain, thereby enhancing their performance in multimodal affective computing tasks. In multimodal learning, cross-modal attention dynamically aligns and focuses on relevant information from different modalities, enhancing the model’s ability to capture sentiment by highlighting key features and their interactions. In multi-task learning, shared representations across affective computing tasks and modalities improve performance by capturing common sentiment-related features from text, audio, and video. More recently, the studies of multimodal learning have advanced the field by pre-training multimodal models on extensive multimodal datasets over years, further improving the performance on downstream tasks such as multimodal sentiment analysis [28, 29, 30, 31]. With the scaling of the pre-trained model, parameter-efficient transfer learning emerges such as adapter [32], prompt [33], instruction-tuning [34] and in-context learning [35, 36].
@@
More and more works of multimodal affective computing leverage these parameter-efficient transfer learning methods to transfer knowledge from pre-trained models (e.g., unimodal pre-trained model or multimodal pre-trained model) to downstream affective tasks to improve model performance by further fine-tuning the pre-trained model. For instance, Zou et al. [37] design a multimodal prompt Transformer (MPT) to perform cross-modal information fusion. UniMSE [38] proposes an adapter-based modal fusion method, which injects acoustic and visual signals into the T5 model to fuse them with multi-level textual information.
@@
Multimodal affective computing encompasses tasks like sentiment analysis, opinion mining, and emotion recognition using modalities such as text, audio, images, video, physiological signals, and haptic feedback. This survey focuses mainly on three key modalities: natural language, visual signals, and vocal signals. We highlight four main tasks in this survey: Multimodal Sentiment Analysis (MSA), Multimodal Emotion Recognition in Conversation (MERC), Multimodal Aspect-Based Sentiment Analysis (MABSA), and Multimodal Multilabel Emotion Recognition (MMER). A considerable volume of studies exists in the field of multimodal affective computing, and several reviews have been published [39, 40, 41, 15, 42, 43]. However, these reviews primarily focus on specific affective computing tasks or specific single modality and overlook an overview of multimodal affective computing across multiple tasks, and the consistencies and differences among these tasks. The goal of this survey is twofold. First, this survey aims to provide a comprehensive overview of multimodal affective computing for beginners exploring deep learning in emotion analysis, detailing tasks, inputs, outputs, and relevant datasets. Second, it also offers insights for researchers to reflect on past developments, explore future trends, and examine technical approaches, challenges, and research directions in areas such as multimodal sentiment analysis and emotion recognition.
@@ IV Modal Feature Extractor
For audio modality, raw acoustic input needs to be processed into numerical sequential vectors. The common way is to use librosa [^2] https://github.com/librosa/librosa. to extract Mel-spectrogram as audio features. It is the short-term power spectrum of sound and is widely used in modern audio processing. Transformer structure has achieved tremendous success of in the field of NLP and computer vision. Gong et al. [64] propose audio spectrogram Transformer (AST), which converted waveform into a sequence of 128-dimensional log Mel filterbank (fbank) features to encode audio modality.
@@ VI-B Pre-trained Model
In recent years, large language model (LLM) [59, 212] and multimodal pre-trained model [213, 27, 22, 214] has achieved significant progress [26, 212, 215]. Compared with non-pretrained model, pre-trained model contains massive transferred knowledge [32, 28], which can be introduced into multimodal representation learning to probe the richer information. Fig. 6 shows the use of pre-trained model in multimodal affective learning task.
@@ VI-C Enhanced Knowledge
External knowledge in machine learning and AI refers to information from outside the training dataset, including knowledge bases, text corpora, knowledge graphs, pre-trained models, and expert insights. Integrating this knowledge can improve performance, generalization, interpretability, and robustness to noisy or limited data [235, 236]. Fig. 7 shows the common way of incorporating external knowledge into multimodal affective learning task.
@@ VI-D4 Multimodal Multi-label Emotion Recognition
MERC MELD [243] T,A,V Friends TV English 13,707
IEMOCAP [244] T,A,V Act English 7,532
@@ VII-B Multimodal Emotion Recognition in Conversation
  • IEMOCAP [244] has 7,532 video clips of multi-party conversations, with labels following Ekman’s six universal emotions, including joy, sadness, fear, anger, surprise and disgust.
@@ IX-A Other Multimodal Affective Computing
Facial expression recognition has significantly evolved over the years, progressing from static to dynamic methods. Initially, static facial expression recognition (SFER) relied on single-frame images, utilizing traditional image processing techniques such as Local Binary Patterns (LBP) and Gabor filters to extract features for classification. The advent of deep learning brought Convolutional Neural Networks (CNNs), which markedly improved the accuracy of SFER [264, 265, 266, 267]. However, static methods were limited in capturing the temporal dynamics of facial expressions [268]. Some methods attempt to approach the problem from a local-global feature perspective, extracting more fine-grained visual representations and identifying key informative segments [269, 270, 271, 272, 273, 274]. These approaches enhance robustness against noisy frames, enabling uncertainty-aware inference. To further enhance accuracy, recent advancements in DFER focus on integrating multimodal data and employing parameter-efficient fine-tuning (PEFT) to adapt large pre-trained models for enhanced performance [275, 276, 277], while Liu et al. [278] introduces the concept of expression reenactment (i.e. normalization), harnessing generative AI to mitigate noise in in-the-wild datasets. Moreover, the burgeoning evidential deep learning (EDL) has shown considerable promise by enabling explicit uncertainty quantification through the distributional measurement in latent spaces for improved interpretability, with demonstrated efficacy in zero-shot learning [279], multi-view classification [280, 281, 282], video understanding [283, 284, 285] and multi-modal named entity recognition.
@@ X Future work
Natural language (spoken and written), visual data (images and videos), and auditory signals (speech, sound, and music) have long been central to multimodal affective computing. Recently, new sensing data types, like haptic and ECG signals, are gaining attention [324]. Haptic signals, which involve touch and convey sensory and emotional attributes, enhance user experiences in areas such as gaming, virtual reality, and mobile apps [325]. These signals offer immediate feedback and can improve user engagement. As research progresses, less-studied modalities like haptic will likely become crucial, complementing established methods and advancing the field of affective computing.
@@ XI Conclusion
Multimodal Affective Computing has emerged as a crucial research direction in artificial intelligence with significant progress in understanding and interpreting emotions. This survey provides a comprehensive overview of the diverse tasks associated with multimodal affective computing, covering its research background, definitions, related work, technical approaches, benchmark dataset, and evaluation metrics. We group multimodal affective computing across MSA, MERC, MABSA and MMER tasks into four categories: multi-task learning, pre-trained modal, enhanced knowledge and context information. Additionally, we summarize the consistency and differences among various affective computing tasks. Also, we report the inherent challenges in multimodal sentiment analysis and explore potential directions for future research and development.